# AraLingBench A Human-Annotated Benchmark for Evaluating Arabic Linguistic Capabilities of Large Language Models

### v1

@@ 1 Introduction
Natural language processing (NLP) in Arabic has seen remarkable progress in recent years, supported by the growing availability of large language models (LLMs) trained specifically for Arabic or Arabic-English bilingual contexts. These models range from general-purpose systems to domain-focused variants (al-khalifa202501340v2; mashaabi202420238v2), and collectively have advanced Arabic NLP in tasks of generation, reasoning, and understanding. However, this rapid growth in model development has not been accompanied by an equally systematic approach to evaluation. The field still lacks reliable methods to assess how well these models truly understand Arabic at the linguistic level.
Existing benchmarks such as BALSAM (sengupta2024balsam), CamelEval (qian2024cameleval), and 3LM (boussaha-etal-2025-3lm) offer a broad coverage of reasoning, comprehension, and task-specific evaluation. However, they share a common limitation: their strong emphasis on factual recall and problem-solving often comes at the expense of testing fundamental language competence. Datasets like EXAMS and ArabicMMLU primarily assess academic or general knowledge, rarely examining whether models can correctly handle grammatical agreement, morphological derivation, or orthographic conventions. Even recent efforts that incorporate dialectal content, such as DialectalArabicMMLU (altakrori202527543v1), remain largely within this knowledge-oriented paradigm. Consequently, the community lacks a benchmark that isolates and measures the linguistic foundations underlying genuine Arabic language understanding.
@@ 2.1 Arabic Language Models
The development of Arabic language models has progressed through several architectural generations. The encoder-based era began with AraBERT (antoun-etal-2020-arabert), which adapted BERT’s architecture for Arabic text, followed by specialized variants such as MARBERT (abdul-mageed-etal-2021-arbert) for dialectal processing, ARBERT (panboonyuen2025albert) for social media content, and CAMeLBERT (mashaabi202420238v2). These models established strong baselines for Arabic NLP tasks but remained limited to discriminative applications such as classification and sequence labeling.
The transition to generative architectures marked a major turning point for Arabic NLP. AraGPT2 introduced Arabic text generation capabilities, paving the way for larger bilingual and multilingual systems. JAIS (sengupta202316149v2) (13B–30B parameters) stands among the most ambitious bilingual Arabic–English efforts, while ALLaM (allam2024allam) (7B–13B) emphasizes instruction tuning and task generalization. Similarly, AceGPT (huang2024acegpt) focuses on cross-lingual transfer learning built upon LLaMA-2, while models such as Hala (hammoud202514008v1) leverage synthetic bilingual data for fine-tuning.
Recent contributions have diversified further. Yehia (yehia2025) has consistently ranked among top-performing Arabic models across multiple benchmarks in our evaluation. Fanar (team2025fanar), developed by QCRI, benefits from extensive computational resources, while domain-specialized systems such as Atlas-Chat (aloui2024atlas) and ArabianGPT target dialectal and regional variants, including Moroccan Darija and Gulf Arabic. Collectively, these developments illustrate the growing maturity and variety of approaches in the Arabic LLM ecosystem.
@@
CAMeLBERT (mashaabi202420238v2) 110M Encoder (BERT) Pretrained from scratch
@@
AceGPT(huang2024acegpt) 7B / 13B Decoder Continued pretraining (Llama-2) + SFT
@@
Fanar (team2025fanar) 9B Decoder Continued pretraining (Gemma-2-9B) + SFT
@@ 2.2 Evaluation Benchmarks for Arabic
In addition, domain-specific benchmarks now target specialized forms of reasoning and linguistic variation. For instance, ArabLegalEval (hassan2024arablegaleval) tests legal reasoning, MedArabiQ (mrini2024medArabiq) evaluates medical knowledge, and several financial datasets assess economic understanding. Cultural and dialectal benchmarks, including Palm (zhang2024palm), AraDiCE (ghanmi2024aradice), and ArabCulture (damak2024arabculture), explore models’ sensitivity to regional variation and sociocultural context.
@@
MedArabiQ(mrini2024medArabiq) 2025 Medical Mixed Mixed No
CamelEval (qian2024cameleval) 2024 Instruction Open Mixed No
AraDiCE (ghanmi2024aradice) 2024 Dialectal + Cultural Mixed Mixed No
@@
ACVA (huang2024acegpt) 2023 Cultural Values MC Mixed No

### v2

@@ 1 Introduction
Natural language processing (NLP) in Arabic has seen remarkable progress in recent years, supported by the growing availability of large language models (LLMs) trained specifically for Arabic or Arabic-English bilingual contexts. These models range from general-purpose systems to domain-focused variants (al-khalifa202501340v2; inoue-etal-2021-interplay), and collectively have advanced Arabic NLP in tasks of generation, reasoning, and understanding. However, this rapid growth in model development has not been accompanied by an equally systematic approach to evaluation. The field still lacks reliable methods to assess how well these models truly understand Arabic at the linguistic level.
Existing benchmarks such as BALSAM (sengupta2024balsam), CamelEval (qian2024camelevaladvancingculturallyaligned), and 3LM (boussaha-etal-2025-3lm) offer a broad coverage of reasoning, comprehension, and task-specific evaluation. However, they share a common limitation: their strong emphasis on factual recall and problem-solving often comes at the expense of testing fundamental language competence. Datasets like EXAMS and ArabicMMLU primarily assess academic or general knowledge, rarely examining whether models can correctly handle grammatical agreement, morphological derivation, or orthographic conventions. Even recent efforts that incorporate dialectal content, such as DialectalArabicMMLU (altakrori202527543v1), remain largely within this knowledge-oriented paradigm. Consequently, the community lacks a benchmark that isolates and measures the linguistic foundations underlying genuine Arabic language understanding.
@@ 2.1 Arabic Language Models
The development of Arabic language models has progressed through several architectural generations. The encoder-based era began with AraBERT (antoun-etal-2020-arabert), which adapted BERT’s architecture for Arabic text, followed by specialized variants such as MARBERT (abdul-mageed-etal-2021-arbert) for dialectal processing, ARBERT (panboonyuen2025albertadvancedlocalizationbidirectional) for social media content, and CAMeLBERT (inoue-etal-2021-interplay). These models established strong baselines for Arabic NLP tasks but remained limited to discriminative applications such as classification and sequence labeling.
The transition to generative architectures marked a major turning point for Arabic NLP. AraGPT2 introduced Arabic text generation capabilities, paving the way for larger bilingual and multilingual systems. JAIS (sengupta202316149v2) (13B–30B parameters) stands among the most ambitious bilingual Arabic–English efforts, while ALLaM (allam2024allam) (7B–13B) emphasizes instruction tuning and task generalization. Similarly, AceGPT (huang-etal-2024-acegpt) focuses on cross-lingual transfer learning built upon LLaMA-2, while models such as Hala (hammoud202514008v1) leverage synthetic bilingual data for fine-tuning.
Recent contributions have diversified further. Yehia (yehia2025) has consistently ranked among top-performing Arabic models across multiple benchmarks in our evaluation. Fanar (fanarteam2025fanararabiccentricmultimodalgenerative), developed by QCRI, benefits from extensive computational resources, while domain-specialized systems such as Atlas-Chat (aloui2024atlas) and ArabianGPT target dialectal and regional variants, including Moroccan Darija and Gulf Arabic. Collectively, these developments illustrate the growing maturity and variety of approaches in the Arabic LLM ecosystem.
@@
CAMeLBERT (inoue-etal-2021-interplay) 110M Encoder (BERT) Pretrained from scratch
@@
AceGPT(huang-etal-2024-acegpt) 7B / 13B Decoder Continued pretraining (Llama-2) + SFT
@@
Fanar (fanarteam2025fanararabiccentricmultimodalgenerative) 9B Decoder Continued pretraining (Gemma-2-9B) + SFT
@@ 2.2 Evaluation Benchmarks for Arabic
In addition, domain-specific benchmarks now target specialized forms of reasoning and linguistic variation. For instance, ArabLegalEval (hassan2024arablegaleval) tests legal reasoning, MedArabiQ (daoud202503427v2) evaluates medical knowledge, and several financial datasets assess economic understanding. Cultural and dialectal benchmarks, including Palm (zhang2024palm), AraDiCE (alshaikh202518442v1), and ArabCulture (sadallah-etal-2025-commonsense), explore models’ sensitivity to regional variation and sociocultural context.
@@
MedArabiQ(daoud202503427v2) 2025 Medical Mixed Mixed No
CamelEval (qian2024camelevaladvancingculturallyaligned) 2024 Instruction Open Mixed No
AraDiCE (alshaikh202518442v1) 2024 Dialectal + Cultural Mixed Mixed No
@@
ACVA (huang-etal-2024-acegpt) 2023 Cultural Values MC Mixed No